Mashed Potato Inc. purchased a machine in January of 2016 for $20,000. The machine had an estimated useful life of 10 years, no salvage value, and was depreciated on a straight-line basis. At the beginning of 2019, Mashed Potato overhauled the machine at a cost of $50,000, which extended the machine’s original life by 9 years. What is the amount of depreciation expense for the machine in 2021?
Answer #1

Accumulated depreciation (2016 to 2018) = 20000/10*3 = 6000

Revised Depreciation = (20000-6000+50000)/16 = 4000 per year

Depreciation expense for the machine in 2021 = $4000